# Latent Reasoning in LLMs as a Vocabulary-Space Superposition

### v1

@@ 4.1 Generating Latent Tokens via Induction-Supervision Masking
Training Schedule.
In practice, the strong masking constraints (LTIM/LTSuM) make naive SFT optimization difficult. We therefore adopt an EM-style alternating schedule. Let the complete set of parameters be $\theta\!=\!(\theta_{\mathrm{enc}},\theta_{\mathrm{llm}})$. First, with the LLM frozen, we optimize the encoder only, $\theta_{\mathrm{enc}}^{t+1}\!=\!\mathrm{arg}\;\mathrm{min}_{\theta_{\mathrm{enc}}}\mathcal{L}_{\mathrm{sup}}(\theta_{\mathrm{enc}},\theta_{\mathrm{llm}}^{t})$. Next, with the encoder frozen, we optimize the LLM, $\theta_{\mathrm{llm}}^{t+1}\!=\!\mathrm{arg}\;\mathrm{min}_{\theta_{\mathrm{llm}}}\mathcal{L}_{\mathrm{sup}}(\theta_{\mathrm{enc}}^{t+1},\theta_{\mathrm{llm}})$. Finally, we jointly fine-tune both components by minimizing, $(\theta_{\mathrm{enc}},\theta_{\mathrm{llm}})\!\leftarrow\!\mathrm{arg}\;\mathrm{min}_{\theta_{\mathrm{llm}},\theta_{\mathrm{enc}}}\mathcal{L}_{\mathrm{sup}}(\theta_{\mathrm{enc}},\theta_{\mathrm{llm}})$. This alternating–joint procedure stabilizes training under the constraints while steadily improving convergence.
4.2 Training the LLM to Autonomously Generate Latent Tokens
After generating latent tokens with the three desired characteristics, we concatenate them with the explicit tokens. The resulting sequence $X\!=\![Q,\text{\textless think\textgreater},z_{1},...,z_{n},\text{\textless/think\textgreater},\mathrm{Answer}]$ then serves as the input for Stage 2 (in Figure 4(b)). Denote the index sets of latent and explicit positions by $S_{\mathrm{lat}}$, $S_{\mathrm{exp}}$, respectively. We use $p_{t}=\alpha_{t}$ as the soft label for the latent position, and ground-truth tokens $y_{t}$ for explicit positions. Let $q_{t}=\mathrm{softmax}(W^{T}h_{t})$ be the decoder’s predictive distribution at position $t$. We optimize the decoder with a KL term on latent slots and a CE term on explicit slots:
$\mathcal{L}_{\mathrm{auto}}(\theta_{\mathrm{llm}})=\lambda\cdot\frac{1}{|S_{\mathrm{lat}}|}\!\sum_{t\in S_{\mathrm{lat}}}\mathrm{KL}\!\big(p_{t}\,|\,q_{t}\big)\;+\;\beta\cdot\frac{1}{|S_{\mathrm{exp}}|}\!\sum_{t\in S_{\mathrm{exp}}}\big(-\log q_{t}[\,y_{t}\,]\big).$ (5)
Optionally, a temperature T>0T\!>\!0 can be applied to both teacher and student distributions (with the usual $T^{2}$ scaling) and $p_{t}$ may be pruned to top-$k$ then renormalized. Minimizing $\mathcal{L}_{\mathrm{auto}}$ trains the decoder to autonomously generate latent token sequences $z_{1},...,z_{N}]$ along with the final answer.
5 Experiments
5.1 Experimental setup
Datasets.
To comprehensively assess the effectiveness of our method, we conducted experiments on both low-difficulty mathematical reasoning datasets (GSM8k-Aug (Deng et al., 2024), GSM-Hard (Gao et al., 2023), SVAMP (Patel et al., 2021), and MultiArith (Roy & Roth, 2015)) and high-difficulty datasets (Math500 (Hendrycks et al., 2021) and AIME24). For the low-difficulty tasks, we trained on the GSM8k-Aug training split to ensure fair comparison with prior work. For the high-difficulty tasks, due to limited computational resources, we used a subset of Open-R1 (Hugging Face, 2025) containing explicit reasoning chains shorter than 4k tokens, resulting in 57,544 training samples. This restriction inevitably constrained the performance of our method.
Baselines and Models.
The baselines considered in our experiments include COT-SFT (Wei et al., 2022), iCOT (Deng et al., 2024), COCONUT (Hao et al., 2024), CODI (reproduced) (Tan et al., 2025), and CoLaR (Tan et al., 2025). For low-difficulty tasks, we adopt LLaMA-3.2-1B-Instruct (Dubey & et al., 2024), while for high-difficulty tasks, we employ DeepSeek-Distill-Qwen-7B (DeepSeek-AI & et al., 2025).
Evaluation Metrics and Details.
Our evaluation measures both accuracy (Pass@N) and efficiency (# L, the number of tokens in the reasoning chain). For the low-difficulty datasets, we conducted five runs with different random seeds. For the high-difficulty datasets, Math500 was evaluated four times, while AIME24 was evaluated 64 times. We report Pass@1 and # L as the mean ± 95% confidence interval (CI). To jointly assess accuracy and efficiency, we also report their ratio, Pass@1/# L. Additional implementation details are provided in Appendix B.
@@ 5.1 Experimental setup
Table 1: Experimental results on four low-difficulty datasets. The number in brackets represents the compression ratio $r$. “ ” marks the best result, and “ ” marks the second best.   - Hidden State refers to removing the Representation Alignment step and directly using the last-layer hidden state as the representation of latent tokens.   - w/o LTIM and   - w/o LTSuM denote removing the Latent Token Induction Mask and Latent Token Supervision Mask, respectively, reducing the model to a standard autoregressive attention mechanism.
@@
Pass@1 # L Pass@1 # L Pass@1 # L Pass@1 # L Pass@1 # L Pass@1/# L
[…]
@@ 5.2 Main Results
Table 1 compares Latent-SFT with state-of-the-art baselines under different compressed ratio $r$ settings. Relative to the latent reasoning baseline, Latent-SFT consistently achieves higher accuracy, shorter reasoning chains, and substantial gains in the Pass@1 metric. These results confirm that defining latent tokens as linear combinations of the vocabulary matrix—thus aligning them with the same semantic space as explicit tokens—is more effective than the traditional last-hidden-state definition. Furthermore, Latent-SFT demonstrates strong out-of-distribution generalization, achieving state-of-the-art performance on all three out-of-distribution datasets and even surpassing COT-SFT on MultiArith. This advantage arises from our latent token definition, which is more compatible with explicit tokens and thus better leverages token representations learned during pre-training, leading to improved generalization.
Table 2: Experimental results on high-difficulty datasets. We primarily examine the performance gap between Soft Embedding and hidden state representations. Long inference chains amplify error accumulation in latent reasoning, while training restricted to lengths of 4k further limits performance. Consequently, the accuracy drop on AIME24 is acceptable. Since prior latent baselines were neither trained nor evaluated on high-difficulty datasets, their results are not reported.
Model MATH-500 AIME24
Pass@1 # L Pass@1 # L
COT-SFT 92.6±1.19 3616±232 54.4±1.74 12992±346
Latent-SFT(Hidden State) 67.8±0.94 698.1±22.1 7.78±2.08 1509±114
Latent-SFT(Soft Embedding) 79.8±1.35 ($12.0\%\uparrow$) 759.3±36.9 19.2±1.60 ($11.4\%\uparrow$) 1484±420
It is noteworthy that Latent-SFT(2) surpasses explicit COT-SFT in Pass@1 on both GSM8k-Aug and MultiArith, while reducing inference chain length by roughly one quarter on average. This finding demonstrates that explicit reasoning does not represent the upper bound of latent reasoning methods. With an appropriate definition and learning strategy for latent tokens, latent reasoning can exceed explicit reasoning models in accuracy. When considering both Pass@1 and #L, Latent-SFT delivers nearly a fourfold improvement over traditional COT-SFT, demonstrating strong overall performance. The real inference example is provided in the Figure 10.
@@
Experimental results on high-difficulty mathematical tasks are shown in Table 2. We found that the accuracy of Latent-SFT on the AIME24 dataset dropped sharply compared to the COT-SFT model. We speculate that this is primarily due to the increased difficulty of learning latent reasoning models with long reasoning chains. Simply put, as reasoning chains length increases, it becomes more difficult to learn the three features required for latent tokens, resulting in a deterioration in latent token representation. This deterioration in latent token representation leads to cumulative errors in the latent reasoning stage, ultimately resulting in a drop in accuracy. Furthermore, due to training resource limitations, our model’s training examples have reasoning chains no longer than 4k in length, further hindering its generalization to the AIME24 dataset, which requires an average of 12k inference chains. (As evidence, our model performs significantly better on the MATH500 dataset, which requires shorter reasoning chains, than on AIME24, which requires longer chains.) Therefore, adapting latent reasoning to high-difficulty mathematical reasoning datasets remains a challenging task.
In terms of inference chain length, Latent-SFT remains effective, achieving substantial reductions on both datasets. More importantly, Soft Embedding, the core definition proposed in this work, consistently outperforms Hidden State even on high-difficulty datasets, thereby validating the generalizability of Definition 1.
@@ 5.3 Ablation Study
To comprehensively assess the effectiveness of Latent-SFT, we perform an ablation study focusing on its three core components for latent token generation: LTIM, LTSuM, and Soft Embedding.
Replacing LTIM and LTSuM with Standard Autoregressive Attention.
LTIM is employed during the information compression process to restrict each special token to attend only to preceding explicit reasoning steps, while preventing attention among special tokens themselves. When replaced with standard autoregressive attention, the hidden states of later special tokens may be influenced by earlier ones. In the early stages of training, before convergence, this effect can be particularly detrimental. As shown in Table 1, across Latent-SFT models with different $r$ values, removing LTIM leads to a substantial drop in Pass@1 (about 5% on average), underscoring the necessity of LTIM.
LTSuM is applied in the supervised decoding process to ensure that each latent token can decode the subsequent explicit reasoning steps, thereby maintaining semantic correctness. Replacing it with standard autoregressive attention—i.e., removing all explicit reasoning sub-segments and allowing only latent tokens to generate the final answer—eliminates intermediate supervision during latent token learning and may cause semantic entanglement. As shown in Table 1, removing LTSuM leads to a substantial drop in accuracy, demonstrating its necessity.
Replace Latent Token Representation from Soft Embedding to the Last Hidden State
Soft Embedding serves as our core definition of latent tokens. While Table 1 shows that it already outperforms traditional last-hidden-state methods such as COCONUT and CODI, for a fair comparison we replace Soft Embedding with the last hidden state in our framework while keeping all other settings unchanged. As reported in Tables 1 and Table 2, Soft Embedding achieves substantially better performance than the last hidden state on both low- and high-difficulty mathematical reasoning tasks, thereby empirically validating its effectiveness.
Figure 5: Latent-SFT’s ECR@K values on GSM8k-Aug test samples. Each curve represents the per-sample ECR@K, with samples sorted along the x-axis by their ECR@300 values.
6 Analysis: Understanding the Essence of Latent Reasoning
In order to probe the fundamental nature of latent reasoning, we raise a central question: What, precisely, does latent reasoning capture during the reasoning process? Inspired by COCONUT’s perspective that latent reasoning explores multiple reasoning paths simultaneously, we propose the following two hypotheses.
Hypothesis 1. Latent Reasoning Represents the Compression of a Single Reasoning Path.

### v2

@@ 4 Latent-Chain: Forming Latent Chain via Induction-Supervision Masking
Training Procedure.
The strict constraints imposed by LTIM and LTSuM can destabilize naive SFT optimization. To address this, we adopt an EM-style alternating optimization schedule. Let $\theta=\{\theta_{\mathrm{enc}},\theta_{\mathrm{llm}}\}$ represent the parameters, where $\theta_{\mathrm{enc}}$ denotes the parameters active during latent token induction. The training proceeds in three phases:
  1. 1.
    Encoder Optimization: Freezing the LLM backbone, we update $\theta_{\mathrm{enc}}$ to minimize reconstruction loss: $\theta_{\mathrm{enc}}^{(t+1)}\leftarrow\mathrm{arg}\min_{\theta_{\mathrm{enc}}}\mathcal{L}_{\mathrm{sup}}(\theta_{\mathrm{enc}},\theta_{\mathrm{llm}}^{(t)})$.
  2. 2.
    LLM Optimization: Freezing the encoder, we update the backbone to adapt to the latent representations: $\theta_{\mathrm{llm}}^{(t+1)}\leftarrow\mathrm{arg}\min_{\theta_{\mathrm{llm}}}\mathcal{L}_{\mathrm{sup}}(\theta_{\mathrm{enc}}^{(t+1)},\theta_{\mathrm{llm}})$.
  3. 3.
    Joint Fine-tuning: Finally, we jointly optimize both components to ensure global convergence.
This progressive strategy stabilizes the training dynamics, allowing the model to gradually adapt to the distributional shift introduced by the latent tokens.
5 Latent-Optim: Training LLMs for Latent Reasoning
Having constructed the Latent-Chain with the requisite semantic properties in Section 4, we address the final question: how can we empower the LLM to generate these high-quality latent chains during reasoning autonomously? Unlike existing methods that rely on inefficient indirect backpropagation from final supervision—thereby lacking explicit guidance for the intermediate latent process—we propose Latent-Optim, a training strategy to learn robust reasoning chains while ensuring generalization.
We concatenate latent chains with the explicit context to form the training sequence $X=[Q,\texttt{<think>},\mathbf{z}_{1},\dots,\mathbf{z}_{N},\texttt{</think>},\mathrm{Answer}]$. A straightforward optimization approach is to minimize the KL divergence between the static latent chain $\bm{\alpha}_{t}$ derived from the encoder and the model’s predicted distribution $\mathbf{q}_{t}=\mathrm{softmax}(\mathbf{W}^{\top}\mathbf{h}_{t})$, combined with the standard cross-entropy loss for explicit tokens:
$\begin{split}\mathcal{L}_{\mathrm{static}}&=\frac{1}{|S_{\mathrm{lat}}|}\sum_{t\in S_{\mathrm{lat}}}\mathrm{KL}(\bm{\alpha}_{t}\,||\,\mathbf{q}_{t})\\ &\quad+\frac{1}{|S_{\mathrm{exp}}|}\sum_{t\in S_{\mathrm{exp}}}\big(-\log q_{t}[y_{t}]\big).\end{split}$ (5)
However, directly supervising the LLM with static latent chains $\bm{\alpha}_{t}$ can lead to overfitting on the specific artifacts of the encoder, thereby limiting the model’s generalization capability.
To address this, we propose a Stochastic Latent Optimization strategy. Instead of treating the latent targets as immutable labels, we view them as parameters of a generative distribution. During each training epoch, we dynamically sample a perturbed target distribution using the Gumbel-Softmax trick (Wu et al., 2025a). Specifically, let $S_{\mathrm{lat}}$ and $S_{\mathrm{exp}}$ denote the index sets for latent and explicit positions, respectively. For a latent position $t\in S_{\mathrm{lat}}$, let $\bm{\alpha}_{t}$ be the static latent chain. We construct a stochastic target $\tilde{\mathbf{p}}_{t}$ by injecting Gumbel noise:
$\tilde{\mathbf{p}}_{t}=\mathrm{softmax}\left(\frac{\log\bm{\alpha}_{t}+\mathbf{g}_{t}}{\tau}\right),\,\mathbf{g}_{t}\sim\mathrm{Gumbel}(0,1),$ (6)
in our experiments, we set the temperature $\tau=1$, which simplifies the unperturbed expectation to the original target $\bm{\alpha}_{t}$ while introducing crucial stochasticity. The noise vector $\mathbf{g}_{t}$ is resampled independently at every training step, ensuring that the model is never exposed to the exact same latent target twice. This dynamic perturbation forces the model to learn the underlying semantic manifold rather than memorizing specific point estimates.
Consequently, the final optimization objective is reformulated as the expectation over this noise:
$\begin{split}\mathcal{L}_{\mathrm{auto}}(\theta_{\mathrm{llm}})&=\frac{1}{|S_{\mathrm{lat}}|}\sum_{t\in S_{\mathrm{lat}}}\mathbb{E}_{\mathbf{g}}\left[\mathrm{KL}\big(\tilde{\mathbf{p}}_{t}(\mathbf{g})\,||\,\mathbf{q}_{t}\big)\right]\\ &\quad+\frac{1}{|S_{\mathrm{exp}}|}\sum_{t\in S_{\mathrm{exp}}}\big(-\log q_{t}[y_{t}]\big).\end{split}$ (7)
By minimizing the expectation over the Gumbel noise, we effectively smooth the optimization landscape. As we formally prove in Theorem 1 (see Appendix C), this stochastic injection acts as an implicit regularizer that penalizes the sharpness of the loss Hessian, thereby guiding the model toward flatter minima and enhancing generalization across diverse reasoning paths.
@@ 5 Latent-Optim: Training LLMs for Latent Reasoning
Table 1: Experimental results on four low-difficulty datasets. The number in brackets represents the compression ratio $r$. “ ” marks the best result, and “ ” marks the second best.   - Hidden State refers to removing the Representation Alignment step and directly using the last-layer hidden state as the representation of latent tokens.   - w/o Gumbel indicates removing the stochastic Gumbel noise injection, effectively performing deterministic optimization on static soft labels.   - w/o LTIM and   - w/o LTSuM denote removing the Latent Token Induction Mask and Latent Token Supervision Mask, respectively, reducing the model to a standard autoregressive attention mechanism.
@@
Pass@1 $\uparrow$ # L $\downarrow$ Pass@1 $\uparrow$ # L $\downarrow$ Pass@1 $\uparrow$ # L $\downarrow$ Pass@1 $\uparrow$ # L $\downarrow$ Pass@1 $\uparrow$ # L $\downarrow$ Pass@1/# L $\uparrow$
LLaMA-3.2-1B-Instruct
CoT-SFT 49.4±.6 70.1±.2 12.1±.2 78.4±.1 60.5±.4 37.2±.1 93.2±.5 38.5±.1 53.8 56.5 0.95
@@
Latent-SFT(4) 46.8±.1 6.81±.0 11.2±.2 7.77±.2 57.6±.3 3.60±.1 94.0±.2 4.15±.0 52.4 5.58 9.39
- w/o Gumbel 44.6±.0 6.64±.0 10.2±.0 7.68±.0 55.6±.1 3.33±.0 93.6±.0 4.11±.0 51.0 5.44 9.37
- Hidden State 36.7±.4 6.61±.1 8.41±.2 7.77±.0 44.4±.2 3.22±.0 90.3±.1 4.06±.0 45.0 5.41 8.32
- w/o LTIM 38.8±.1 6.59±.1 8.67±.1 8.01±.1 48.9±.1 3.34±.1 90.8±.0 4.36±.0 46.8 5.58 8.39
- w/o LTSuM 42.6±.0 6.70±.0 9.98±.1 7.60±.0 53.9±.1 3.36±.0 92.6±.0 4.27±.0 49.7 5.48 9.07
@@
- w/o Gumbel 50.4±.1 12.4±.0 10.9±.0 13.9±.0 57.8±.1 5.98±.0 93.8±.0 7.21±.0 53.2 9.87 5.39
- Hidden State 41.1±.3 12.8±.0 9.01±.1 14.2±.1 49.6±.2 6.17±.0 91.1±.0 7.29±.0 47.7 10.1 4.72
- w/o LTIM 41.3±.1 12.9±.1 9.32±.0 14.0±.0 53.6±.1 6.25±.1 91.5±.0 7.96±.0 48.9 10.3 4.75
- w/o LTSuM 49.0±.0 12.1±.0 10.3±.0 13.7±.0 56.5±.2 6.22±.0 93.0±.0 7.23±.0 52.2 9.81 5.32
LLaMA-3.2-3B-Instruct
CoT-SFT 61.1±.3 70.5±.2 18.8±.2 80.6±.1 70.9±.2 34.2±.1 98.3±.5 37.4±.1 62.2 55.7 1.12
@@
6 Experiments
6.1 Experimental Setups
Datasets.
To comprehensively assess the effectiveness of our method, we conducted experiments on both low-difficulty mathematical reasoning datasets (GSM8k-Aug (Deng et al., 2024), GSM-Hard (Gao et al., 2023), SVAMP (Patel et al., 2021), and MultiArith (Roy and Roth, 2015)) and high-difficulty datasets (Math500 (Hendrycks et al., 2021) and AIME24). For the low-difficulty tasks, we trained on the GSM8k-Aug training split to ensure fair comparison with prior work. For the high-difficulty tasks, due to limited computational resources, we used a subset of Open-R1 (Hugging Face, 2025) containing explicit reasoning chains shorter than 4k tokens, resulting in 57,544 training samples. This restriction inevitably constrained the performance of our method.
Baselines and Models.
The baselines considered in our experiments include COT-SFT (Wei et al., 2022), iCOT (Deng et al., 2024), COCONUT (Hao et al., 2024), CODI (reproduced) (Tan et al., 2025), and CoLaR (Tan et al., 2025). To align model capabilities with task complexity, we strictly match base models to dataset difficulty. For low-difficulty tasks, we adopt LLaMA-3.2-1B-Instruct and LLaMA-3.2-3B-Instruct (Dubey and et al., 2024), while for high-difficulty tasks, we employ DeepSeek-Distill-Qwen-7B (DeepSeek-AI, 2025; Yang et al., 2025).
Evaluation Metrics and Details.
Our evaluation measures both accuracy (Pass@N) and efficiency (# L, the number of tokens in the reasoning chain). For the low-difficulty datasets, we conducted five runs with different random seeds. For the high-difficulty datasets, Math500 was evaluated four times, while AIME24 was evaluated 64 times. We report Pass@1 and # L as the mean ± 95% confidence interval (CI). To jointly assess accuracy and efficiency, we also report their ratio, Pass@1/# L.
Implementation Details
Since Latent-SFT requires weighting with respect to the vocabulary matrix, we freeze the vocabulary parameters during training. All models are trained using LoRA with rank 64. Following COCONUT, we initialize all methods from the COT-SFT model to accelerate convergence. In practice, we find that a fixed-length segmentation function performs better than a semantic-based segmentation. Therefore, all results reported in this paper are based on fixed-length segmentation functions. For LLaMA-3.2-1B-Instruct and LLaMA-3.2-3B-Instruct, we set the learning rate to $1\times 10^{-4}$for Latent-Chain generation and $4\times 10^{-4}$ for Latent-Optim. For DeepSeek-Distill-Qwen-7B, the learning rates are $3\times 10^{-5}$ (Latent-Chain) and $1\times 10^{-5}$ (Latent-Optim). Following (DeepSeek-AI, 2025), we use a decoding temperature of 0.6 and top-p of 0.95.
During evaluation, low-difficulty mathematical reasoning datasets were inferred using the Transformers library with a maximum sequence length of 128. High-difficulty datasets were evaluated with the improved SGlang library, supporting a maximum token length of 32,768. All training and inference experiments were conducted on 8×A100 GPUs.
Table 2: Experimental results of DeepSeek-Distill-Qwen-7B on high-difficulty datasets. We primarily examine the performance gap between Superposition Embedding and hidden state representations. To ensure a fair comparison, we set the maximum reasoning chain length to 4K tokens for all models.
Model MATH-500 AIME24
@@ 6.2 Main Results
Table 1 compares Latent-SFT with state-of-the-art baselines under different compressed ratio $r$ settings. Relative to the latent reasoning baseline, Latent-SFT consistently achieves higher accuracy, shorter reasoning chains, and substantial gains in the Pass@1 metric. These results confirm that defining latent tokens as a superposition of vocab-space—thus aligning them with the same semantic space as explicit tokens—is more effective than the traditional last-hidden-state definition. Furthermore, Latent-SFT demonstrates strong out-of-distribution generalization, achieving state-of-the-art performance on all three out-of-distribution datasets.
[…]
@@
[…]
@@ 6.3 Ablation Study
Necessity of Masking Mechanisms.
LTIM enforces causal independence to stabilize training, while LTSuM ensures latent tokens retain sufficient semantic information for decoding. As shown in Table 1, removing either mechanism leads to a significant accuracy drop ($\sim$5%), confirming their critical role in maintaining the structural integrity of the reasoning chain.
Latent Representation and Stochastic Optimization.
We further validate our specific design choices. First, replacing Superposition Embedding with raw hidden states degrades performance, demonstrating the importance of aligning latent tokens with the intrinsic vocabulary manifold. Second, ablating Gumbel-noise injection (w/o Gumbel) leads to a noticeable decline, confirming that stochasticity acts as a vital regularizer (Theorem 1) to prevent overfitting and enhance generalization.
Figure 5: ECR@K analysis of Latent-SFT on the GSM8k-Aug test set. Each curve represents the per-sample ECR@K, with samples sorted along the x-axis by their ECR@300 values. The dashed line indicates the baseline for uncompressed single-chain reasoning. Latent-SFT consistently surpasses this baseline, with higher training compression ratios yielding greater information compression. This demonstrates that Latent-SFT effectively compresses single explicit reasoning paths during the latent reasoning.
Figure 6: $N_{\mathrm{eff}}$ values of Latent-SFT on the Multi-Chain GSM8k-Aug dataset. The legend reports the mean values of both $N_{\mathrm{eff}}$ and the Top-2 Score. Overall, Latent-SFT models across various compression ratios consistently exceed the significant parallel threshold. This confirms that the latent reasoning process effectively operates as a superposition of multiple simultaneous reasoning chains.
7 Essence of Latent Reasoning
A central premise of our proposed Latent-SFT is that it enables the model to transcend simple sequence compression and instead operate in a state of reasoning superposition. This raises a fundamental validation question: Has our method truly realized this ideal state of latent reasoning? Specifically, does the learned latent process effectively capture and explore multiple reasoning paths simultaneously, rather than merely memorizing a single trajectory? To rigorously answer this, we investigate the underlying mechanism of Latent-SFT through two hypotheses.
Hypothesis 1. Latent Reasoning as the Compression of a Single Reasoning Chain.